OpenStudy (anonymous):

For a certain bathtub, it takes the hot water faucet 3 times as long to fill the tub as it does the cold water faucet. Left on together, the two faucets take 6 minutes to fill the tub. How long will it take the hot water faucet to fill the tub by itself?

OpenStudy (anonymous):

9 minutes?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Let c= the flow rate for cold water (tub/min) We are given that the flow rate for hot water is 1/3 of the cold water rate. So we have this equation. 1/3c + c = 1tub/6 minutes 4/3c = 1/6 c= 1/6*3/4 c=3/24 c=1/8 or 1 tub/8 minutes So the cold water faucet can fill the tub by its self in 8 minutes. The hot water flow rate is 1/3 of the cold water rate. So 1/3 of 1/8 equals 1/24 or 1 tub/24 minutes So the hot water faucet can fill the tub by its self in 24 minutes. Check: (1tub/24minutes) (6 minutes) + (1 tub/8minutes)(6 minutes) = 6/24 tub + 6/8 tub 6/24 tub + 18/24 tub = 24/24 tub or 1 tub